Stacey would like to have​ $1 million available to her at retirement. Her investments have an average annual return of 12​%. If she makes contributions of ​$130 per​ month, will she reach her goal when she retires in 29 ​years? The amount Stacy has after 29 years would be ​$ nothing. ​(Use your financial calculator and round to the nearest​ cent.)   If she makes contributions of ​$130 per​ month, will she reach her goal when she retires in 29 ​years?
Answer #1

A=130, i=12% per year = 1% every month, i = 0.01,

no. of periods = 29 yrs*12 months = 348

Future value of annuity = A * ( ( 1+i)n -1) / i

Future Value = 130 * {(1+0.01)348 -1) / 0.01 = 130 * 3090.348 = 401745.3

Therefore Stacey would not have 1 million at the end of 29 years.